# End-to-End Audio-Visual Learning for Cochlear Implant Sound Coding in Noisy Environments

### v1

@@ Abstract
The cochlear implant (CI) is a remarkable biomedical device that successfully enables individuals with severe-to-profound hearing loss to perceive sound by converting speech into electrical stimulation signals. Despite advancements in the performance of recent CI systems, speech comprehension in noisy or reverberant conditions remains a challenge. Recent and ongoing developments in deep learning reveal promising opportunities for enhancing CI sound coding capabilities, not only through replicating traditional signal processing methods with neural networks, but also through integrating visual cues as auxiliary data for multimodal speech processing. Therefore, this paper introduces a novel noise-suppressing CI system, AVSE-ECS, which utilizes an audio-visual speech enhancement (AVSE) model as a pre-processing module for the deep-learning-based ElectrodeNet-CS (ECS) sound coding strategy. Specifically, a joint training approach is applied to model AVSE-ECS, an end-to-end CI system. Experimental results indicate that the proposed method outperforms the previous ECS strategy in noisy conditions, with improved objective speech intelligibility scores. The methods and findings in this study demonstrate the feasibility and potential of using deep learning to integrate the AVSE module into an end-to-end CI system.
Index Terms:
cochlear implant (CI), audio-visual speech enhancement (AVSE), joint training, end-to-end model
Figure 1: The architecture of the proposed cochlear implant (CI) system. The area enclosed by the dotted line indicates the AVSE-ECS network.
Figure 2: Overall architecture of the training AVSE-ECS.
@@ I Introduction
The cochlear implant (CI) is a hearing device designed to help individuals with severe-to-profound hearing loss regain partial hearing. The CI system involves an external, upgradable sound processor that uses radio frequency to wirelessly link to a surgically implanted internal device. The CI system transforms speech signals into electrical pulse patterns, which are transmitted via the auditory nervous system to the auditory cortex of the brain, enabling sound perception. Developed through multidisciplinary collaboration, modern CIs now provide impressive speech recognition performance. With well-designed sound coding strategies that translate acoustic signals into electrode stimulation patterns, such as the widely used advanced combination encoder (ACE) strategy [1], most CI users can understand more than 80–90% of speech sentences in quiet conditions [2]. However, further improvements are still needed, particularly in the processing of noisy speech.
@@
Speech enhancement (SE) offers a potential solution for improving CI in noisy environments. With the rapid development of deep learning, SE approaches for CIs are gradually shifting from conventional signal processing algorithms to data-driven neural networks [3, 4, 5]. Many SE approaches have been applied to CIs, including recent end-to-end architectures [6, 7] that integrate both the pre-processing stage and the sound coding strategy [8, 9], the core component responsible for converting audio into electrical stimuli. However, these approaches, which rely solely on audio input, do not make use of the visual cues available for processing background noise and overlapping speech. In general audio applications, audio-visual speech enhancement (AVSE), a common multimodal signal processing approach, has been proposed to incorporate visual information such as lip movements and facial expressions alongside audio signals, and has proven effective in outperforming audio-only speech enhancement methods [10, 11, 12, 13, 14, 15, 16]. These findings encourage further exploration of visual cues and AVSE models for CI sound processing in noisy environments [17, 18].
@@
Improving conventional CI systems with an AVSE module requires careful design to overcome associated challenges. The core signal processing functions of the ACE coding strategy, envelope detection and channel selection (CS), are not differentiable, making them difficult to integrate with neural-network-based AVSE modules and refine through data-driven regression. To overcome this limitation, the ElectrodeNet-CS (ECS) model introduced in [19] is used to explore the capabilities of deep-learning-based CI coding strategies. This model replicates the envelope detection and channel selection processes of the ACE strategy using a deep neural network (DNN) along with a dedicated CS function, while achieving comparable performance to ACE. The ECS coding strategy enables the integration of different front-end and post-processing components, allowing for joint training and end-to-end optimization across various tasks. Consequently, the end-to-end AVSE-ECS CI system, which involves the pre-processing AVSE module and the ECS strategy, is proposed and validated using objective evaluations.
@@
The structure of this paper is as follows. Section II introduces the proposed AVSE-based CI system. Section III describes the experimental setup for objective evaluation. Section IV presents the results. Section V provides the conclusions.
@@ II-A System Architecture
The system architecture of the proposed AVSE-ECS CI system is illustrated in Fig. 1. The AVSE model incorporates auxiliary visual cues to generate enhanced speech, which is further processed by the ECS strategy into electrode stimulation patterns. A tone vocoder [20] is used to convert electrode stimulation patterns into audible speech, which can then be analyzed using objective evaluation methods. Furthermore, this study introduces a joint training approach, and the architecture for training the AVSE-ECS network is illustrated in Fig. 2.
II-B Coding Strategy
@@
This study adopts the DNN-based ECS model proposed in [19] as the coding strategy, which is the core sound processing of the CI system. This model consists of four dense layers with 1024, 512, 256, and 22 neurons, followed by a custom topk layer designed for the CS function. The ECS model is pre-trained before being integrated into the AVSE-ECS system. Pre-training involved paired spectral–temporal matrices across L = 65 bins and M = 22 channels derived from clean speech processed with the ACE strategy. The training was configured to preserve key information in the $N_{topk}=8$ channels selected by the network, while distributing lower envelope data across the remaining $M-N_{topk}$ channels, using the L1 norm across all M channels.
II-C Visual Encoder
@@
The temporal convolution network (TCN) is employed as the visual encoder as illustrated in Fig. 1 and Fig. 2. This model consists of a 3D convolutional layer to process temporal sequences of frames, followed by a 2D ResNet-18. The extracted visual features are then passed to the AVSE model as visual embeddings. Note that a modified ResNet-18 visual encoder is pre-trained using the Lip Reading in the Wild (LRW) dataset [21]. To provide the most relevant visual cues for SE, the encoder focuses on the mouth, referred to as the region of interest (ROI), rather than utilizing the entire facial region. Facial landmarks are detected using the Mediapipe face tracker [22], and the mouth ROI is extracted accordingly. To reduce computational overhead during training, the visual encoder is kept frozen and the embeddings are pre-extracted, minimizing both memory usage and the number of training parameters.
II-D AVSE Pre-processing
@@
For the AVSE pre-processing, this study adopts and modifies the NCSN++ model from [23]. As this original UNet is designed for generative tasks, the timestep is set to 1 in this study, and the model’s output directly serves as the enhanced speech. For audio feature preprocessing before AVSE, the experimental settings are similar to [24]. Audio signals are resampled to 16 kHz and converted into complex-valued STFT representations. Using a window size of 510 and a periodic Hann window, the resulting frequency dimension is F = 256. The hop length is set to 128. Each STFT spectrogram is truncated to T = 256 time frames, enabling random cropping during batch training.
II-E Audio-Visual Fusion
@@
This study explores the use of cross-attention blocks, illustrated by the red block in Fig. 2, to integrate audio and visual information. The attention mechanism enables neural networks to dynamically focus on the most relevant parts of the input, capture contextual dependencies, and facilitate alignment across modalities. Following a similar approach to [25], the self-attention blocks within the UNet architecture are modified by replacing the audio keys and values with visual embeddings generated by the visual encoder. Augmenting the audio input with lip movement information is intended to utilize visual cues to enhance speech reconstruction and thereby improve the overall performance of the SE system. To explain the design of the audio-visual fusion module, the cross-attention mechanism is shown in Fig. 3. Audio features (in blue) and visual features (in green) are provided as input, with dimensions indicated numerically. $\bf M_{Q}$}, $\bf M_{K}$}, and $\bf M_{V}$} represent the transformation layers for query, key, and value, respectively. They are fused by the cross attention block, then reshaped and output as audio.
Figure 3: The schematic diagram of the audio-visual fusion block.
II-F Joint Training
@@
Joint training in this study simultaneously optimizes both the AVSE pre-processing module, represented by the blue block in Fig. 2, and the ECS module, shown as the yellow block in Fig. 2, within an end-to-end framework. To guide the SE module toward generating high-quality enhanced speech during the learning process, a training objective called spectrogram loss is defined as follows:
@@
where $\bf E$} is the enhanced speech from the AVSE model and $\bf C$} is the clean ground truth. To consider the distance across the enhanced electrodogram and the clean electrodogram, the electrodogram loss is defined as
@@ III Experiments
III-A Objective Evaluation Metrics
To assess the performance of CI systems, three commonly-used objective evaluation metrics are employed: short-time objective intelligibility (STOI) [26], extended short-time objective intelligibility (ESTOI) [27], and normalized covariance measure (NCM) [28, 29]. These evaluation metrics are computed based on the correlations between CI vocoded speech and clean speech to estimate speech intelligibility.
III-B Dataset
The audiovisual dataset used in this study is based on the Taiwan Mandarin Hearing in Noise Test (TMHINT) speech material [30], which consists of 16 phonemically balanced lists. Each list contains 20 sentences, with each sentence consisting of 10 syllables (Chinese characters). The duration of each utterance ranges from approximately 2 to 4 seconds. The audio-only dataset and the audio-visual dataset are both based on TMHINT sentences. The audio-only TMHINT dataset is used to train the ECS model, as described in [19], with the exception that 320 clean sentences from Lists 1 to 8 are used in this study.
For training and inference of the AVSE-ECS system, the Taiwan Mandarin Speech with Video (TMSV) dataset [11], consisting of speech and video recordings of TMHINT sentences, is adopted. The method for using the TMSV dataset is similar to that described in [11]. This dataset includes speech recordings from 18 native Mandarin speakers (13 males and 5 females). In this study, considering gender balance, 8 speakers (4 male, 4 female) from TMSV are selected for training and validation. For each speaker with 320 utterances, sentences 1-190 are used for training, and sentences 191-200 for validation. These speech files are mixed with 100 types of noise [31] at 5 signal-to-noise ratio (SNR) levels (-12, -6, 0, 6, 12 dB). To reduce training time, a total of 12,000 utterances are randomly sampled. For evaluation, utterances 201–320 from two unseen speakers (1 male, 1 female) are used. These are corrupted with five noise types including music, street noise, engine noise, baby cry, and pink noise, and mixed at SNR levels of -1, -4, -7, and -10 dB, yielding a total of 4,800 noisy utterances. The training and testing sets are fully mismatched in terms of speech content, noise types, speakers, and SNR levels to ensure robust generalization.
III-C Training Configuration
@@ III-D Objective Evaluation
In the first experiment, an objective evaluation is conducted using the pre-trained ECS model with noisy and clean speech from the TMSV test set. This experiment examines the STOI, ESTOI, and NCM scores of the ECS model under clean and noisy conditions.
The second experiment is designed to investigate the effects of model weights during joint training on speech intelligibility. Following the pre-trained method, $\alpha$ is set to 1, and different values of $\beta$ are explored in this experiment. The results of this experiment are used to select $\beta$ for the next experiment.
The third experiment compares different CI network implementations, including ACE, ECS, ASE-ECS, and AVSE-ECS (Fig. 4).
Figure 4: The schematic comparison between different neural-network-based implementations: (a) ECS, (b) ASE-ECS), and (c) AVSE-ECS. The yellow block stands for ECS, the blue block represents the NCSN++ model, and the green block indicates the TCN visual encoder in Fig. 2.
The ACE strategy is one of the most common approach based on traditional signal processing. It serves as the baseline for performance comparison. The ASE-ECS model is an audio-only system in which self-attention replaces the cross-attention mechanism used in the AVSE-ECS model. AVSE-ECS, the proposed end-to-end CI system, is evaluated using both the pre-trained method and the joint training method. For all experiments, $\alpha$} is set to 1, and in the joint training setup, $\beta$} is set to 0.5.
IV Results
This section presents the results of the objective evaluation.
TABLE I: Objective evaluation of the ECS model
@@ IV Results
TABLE II: Objective evaluation of the AVSE-ECS network
with different $\beta$ values
@@
TABLE III: Objective evaluation of different CI implementations

### v2

# End-to-End Audio-Visual Learning for Cochlear Implant Sound Coding Simulations in Noisy Environments

@@ Abstract
The cochlear implant (CI) is a successful biomedical device that enables individuals with severe-to-profound hearing loss to perceive sound through electrical stimulation, yet listening in noise remains challenging. Recent deep learning advances offer promising potential for CI sound coding by integrating visual cues. In this study, an audio-visual speech enhancement (AVSE) module is integrated with the ElectrodeNet-CS (ECS) model to form the end-to-end CI system, AVSE-ECS. Simulations show that the AVSE-ECS system with joint training achieves high objective speech intelligibility and improves the signal-to-error ratio (SER) by 7.4666 dB compared to the advanced combination encoder (ACE) strategy. These findings underscore the potential of AVSE-based CI sound coding.
preprint: Author, JASA-EL
@@ I Introduction
The cochlear implant (CI) is a hearing device designed to help individuals with severe-to-profound hearing loss regain partial hearing [CI_review_zeng2022]. The CI system involves an external, upgradable sound processor that uses radio frequency to wirelessly link to a surgically implanted internal device. The CI system transforms speech signals into electrical pulse patterns, which are transmitted via the auditory nervous system to the auditory cortex of the brain, enabling sound perception. Developed through multidisciplinary collaboration, modern CIs now provide impressive speech recognition performance. With well-designed sound coding strategies that translate acoustic signals into electrode stimulation patterns, such as the widely used advanced combination encoder (ACE) strategy [ACE_original_n_of_m_vandali2000], most CI users can understand more than 80–90% of speech sentences in quiet conditions [CI_review_zeng2008]. However, further improvements are still needed, particularly in the processing of noisy speech.
@@
Speech enhancement (SE) provides a potential pathway for improving CI in noisy environments. With the rapid development of deep learning, SE approaches for CIs are gradually shifting from conventional signal processing algorithms to data-driven neural networks [BioASP_DDAE_lai2017, NR_CI_henry2023, DCCTN_mamun2024]. Many SE approaches have been applied to CIs, including recent end-to-end architectures [DeepACE_gajecki2023, AdversarialDeepACE_gajecki2025] that integrate both the pre-processing stage and the sound coding strategy [strategy_review_wouters2015, NCU_CI_TNSRE_huang2021], the core component responsible for converting audio into electrical stimuli. However, these audio-only SE (ASE) approaches, which rely solely on audio input, do not make use of the visual cues available for processing background noise and overlapping speech. In general audio applications, audio-visual speech enhancement (AVSE), a common multimodal signal processing approach, has been proposed to incorporate visual information such as lip movements and facial expressions alongside audio signals, and has proven effective in outperforming audio-only speech enhancement methods [AVSE_light_chuang2020, AVSE_light_chuang2022, AVSE1, AVSE2, AVSE3, AVSE4, AVSE5]. Therefore, the exploration of visual cues and AVSE models for CI sound processing in noisy environments remains an important area of ongoing research [BioASP_visualcues_tseng2021, BioASP_AVSE_lai2025].
@@
A potential way to improve the CI system using an AVSE module is to integrate it with the ElectrodeNet-CS (ECS) model [ElectrodeNet_huang2024]. The core signal processing functions of the ACE coding strategy, envelope detection and channel selection (CS), are not differentiable, making them difficult to integrate with neural-network-based AVSE modules and refine through data-driven regression. To overcome this limitation, the ECS model replicates the core signal processing of ACE using a deep neural network (DNN) with a dedicated CS function, achieving comparable performance. Furthermore, the ECS coding strategy enables integration with diverse front-end and post-processing components, as well as joint training and end-to-end optimization. Consequently, the end-to-end AVSE-ECS CI system, which incorporates the AVSE pre-processing module and the ECS model, is proposed and validated using objective evaluations in this study.
@@
This paper is organized as follows. Section II explains the proposed AVSE-based CI system. Section III describes the experimental setup for objective evaluation. Section IV offers the results and discussion. Section V draws the conclusion.
@@ II.1 System Architecture
The system architecture of the proposed AVSE-ECS CI system is illustrated in Fig. 1(a). The AVSE model incorporates auxiliary visual cues to generate enhanced speech, which is further processed by the ECS strategy into electrode stimulation patterns. A tone vocoder [vocoder_tone_dorman2002] is used to convert electrode stimulation patterns into audible speech, which can then be analyzed using objective evaluation methods. Furthermore, this study introduces a joint training approach, and the architecture for training the AVSE-ECS network is illustrated in Fig. 1(b). In this study, three CI network implementations are investigated: ECS (Fig. 2(a)), ASE-ECS (Fig. 2(b)), and AVSE-ECS (Fig. 2(c)).
II.2 Coding Strategy
@@
This study adopts the DNN-based ECS model proposed in [ElectrodeNet_huang2024] as the coding strategy, which is the core sound processing of the CI system. This model consists of four dense layers with 1024, 512, 256, and 22 neurons, followed by a custom topk layer designed for the CS function. The ECS model is pre-trained before being integrated into the AVSE-ECS system. Pre-training involved paired spectral–temporal matrices across L = 65 bins and M = 22 channels derived from clean speech processed with the ACE strategy. The training was configured to preserve key information in the $N_{topk}=8$ channels selected by the network, while distributing lower envelope data across the remaining $M-N_{topk}$ channels, using the L1 norm across all M channels.
Figure 1: (a) The architecture of the proposed cochlear implant (CI) system, and the area enclosed by the dashed rectangle indicates the AVSE-ECS network. (b) Overall joint-training architecture for the AVSE-ECS network.
II.3 Visual Encoder
@@
The temporal convolution network (TCN) is employed as the visual encoder as illustrated in Fig. 1. This model consists of a 3D convolutional layer to process temporal sequences of frames, followed by a 2D ResNet-18. The extracted visual features are then passed to the AVSE model as visual embeddings. Note that a modified ResNet-18 visual encoder is pre-trained using the Lip Reading in the Wild (LRW) dataset [LRW]. To provide the most relevant visual cues for SE, the encoder focuses on the mouth, referred to as the region of interest (ROI), rather than utilizing the entire facial region. Facial landmarks are detected using the Mediapipe face tracker [Mediapipe], and the mouth ROI is extracted accordingly. To reduce computational overhead during training, the visual encoder is kept frozen and the embeddings are pre-extracted, minimizing both memory usage and the number of training parameters.
II.4 AVSE Pre-processing
@@
For the AVSE pre-processing, this study adopts and modifies the NCSN++ model from [ScoreDiff]. As this original UNet is designed for generative tasks, the timestep is set to 1 in this study, and the model’s output directly serves as the enhanced speech. For audio feature preprocessing before AVSE, the experimental settings are similar to [CMKT]. Audio signals are resampled to 16 kHz and converted into complex-valued STFT representations. Using a window size of 510 and a periodic Hann window, the resulting frequency dimension is F = 256. The hop length is set to 128. Each STFT spectrogram is truncated to T = 256 time frames, enabling random cropping during batch training.
II.5 Audio-Visual Fusion
@@
This study explores the use of cross-attention blocks (red block in Fig. 1 (b)) to integrate audio and visual information. The attention mechanism enables neural networks to focus on the most relevant input parts, capture contextual dependencies, and facilitate alignment across modalities. Following a similar approach to [DAVSE], the self-attention blocks within the UNet architecture are modified by replacing the audio keys and values with visual embeddings generated by the visual encoder. Augmenting the audio input with lip movement information is intended to utilize visual cues to enhance speech reconstruction and thereby improve the overall performance of the SE system. To explain the design of the audio-visual fusion module, the cross-attention mechanism is shown in Fig. 2 (d). Audio features (in blue) and visual features (in green) are provided as input, with dimensions indicated numerically. $\bf M_{Q}$}, $\bf M_{K}$}, and $\bf M_{V}$} represent the transformation layers for query, key, and value, respectively. They are fused by the cross attention block, then reshaped as audio output.
Figure 2: Neural network implementations: (a) ECS, (b) ASE-ECS, (c) AVSE-ECS, and (d) cross-attention block. As illustrated in Fig. 1, the color blocks represent the modules of ECS (yellow), NCSN++ (blue), and TCN visual encoder (green).
II.6 Joint Training
@@
Joint training, as shown in Fig. 1 (b), simultaneously optimizes both the AVSE pre-processing module (blue block) and the ECS module (yellow block) within an end-to-end framework. To guide the SE module toward generating high-quality enhanced speech during the learning process, a training objective called spectrogram loss is defined as follows:
@@
where MSE is the mean squared error between $\bf E$}, the enhanced speech from the AVSE model, and $\bf C$}, the clean ground truth. To measure the distance for each enhanced-clean electrodogram pair, the electrodogram loss is defined as
@@ III Experiments
III.1 Objective Evaluation Metrics
To assess the performance of CI systems, three commonly-used objective evaluation metrics are employed: short-time objective intelligibility (STOI) [STOI], extended short-time objective intelligibility (ESTOI) [ESTOI], and normalized covariance measure (NCM) [NCM_2]. These metrics estimate speech intelligibility by measuring the correlation between CI-vocoded and clean speech. Furthermore, the neural models are compared using the signal-to-error ratio (SER) [SER_2015], where the error components include remaining noise after SE processing as well as artifacts and distortions introduced by the SE model.
III.2 Dataset
The audiovisual dataset used in this study is based on the Taiwan Mandarin Hearing in Noise Test (TMHINT) speech material [MHINT_wong2007], which consists of 16 phonemically balanced lists. Each list contains 20 sentences, with each sentence consisting of 10 syllables (Chinese characters). The duration of each utterance ranges from approximately 2 to 4 seconds. The audio-only dataset and the audio-visual dataset are both based on TMHINT sentences. The audio-only TMHINT dataset is used to train the ECS model, as described in [ElectrodeNet_huang2024], with the exception that 320 clean sentences from Lists 1 to 8 are used in this study.
For training and inference of the AVSE-ECS system, the Taiwan Mandarin Speech with Video (TMSV) dataset [AVSE_light_chuang2022], consisting of speech and video recordings of TMHINT sentences, is adopted. The method for using the TMSV dataset is similar to that described in [AVSE_light_chuang2022]. This dataset includes speech recordings from 18 native Mandarin speakers (13 males and 5 females). In this study, considering gender balance, 8 speakers (4 male, 4 female) from TMSV are selected for training and validation. For each speaker with 320 utterances, sentences 1-190 are used for training, and sentences 191-200 for validation. These speech files are mixed with 100 types of noise [Training_Noise] at five signal-to-noise ratio (SNR) levels (-12, -6, 0, 6, 12 dB). To reduce training time, a total of 12,000 utterances are randomly sampled. For evaluation, utterances 201–320 from two unseen speakers (1 male, 1 female) are used. These are corrupted with five noise types including music, street noise, engine noise, baby cry, and pink noise, and mixed at SNR levels of -1, -4, -7, and -10 dB, yielding a total of 4,800 noisy utterances. The training and testing sets are fully mismatched in terms of speech content, noise types, speakers, and SNR levels to ensure robust generalization.
III.3 Training Configuration
@@ III.4 Objective Evaluation
In the first experiment, the pre-trained ECS model is evaluated on the TMSV test set using STOI, ESTOI, and NCM metrics under both clean and noisy conditions. All evaluations are conducted on unseen speakers, with the objective scores for the noisy condition averaged over different noise types and SNR levels.
The second experiment examines how model weights in joint training affect speech intelligibility. Following the pre-trained method with $\alpha$ = 1, and different $\beta$ values are studied. The result helps select $\beta$ for the next experiment.
The third experiment compares three CI network implementations: ECS (Fig. 2(a)), ASE-ECS (Fig. 2(b)), and AVSE-ECS (Fig. 2(c)), with the common ACE coding strategy as the baseline. Apart from the ECS model without the SE function, the ASE-ECS model is an audio-only system in which self-attention replaces the cross-attention mechanism used in the AVSE-ECS model. AVSE-ECS, the proposed end-to-end CI system, is evaluated using both the pre-trained method and the joint training method. For all experiments, $\alpha$} is set to 1, and in the joint training setup, $\beta$} is set to 0.5.
IV Results and Discussion
IV.1 ECS with clean and noisy speech
The overall objective intelligibility of the ECS model under clean and noisy conditions is shown in Table 1. For clean speech, ECS achieves an STOI score of 0.7604, which drops to 0.4870 under noisy conditions, averaged across speakers, noise types and SNR levels. Similar declines in ESTOI and NCM occur as conditions shift from clean to noisy. The ECS model’s performance declines under noisy conditions, as it only simulates the ACE strategy without SE capability.
Table 1: Objective evaluation of the ECS model
@@ IV.1 ECS with clean and noisy speech
Table 2: Objective evaluation of the AVSE-ECS network with different $\beta$ values under noisy conditions
@@
IV.2 AVSE-ECS with different $\beta$ values
Table 2 presents the objective evaluation results of AVSE-ECS with different joint training setups. In noisy conditions, AVSE-ECS attains STOI scores of 0.6295 ($\beta$ = 0.25), 0.6305 ($\beta$ = 0.5), and 0.6196 ($\beta$ = 1), averaged over speakers, noise types and SNR levels. Similar trends in STOI and ESTOI scores appear across the three $\beta$ values. These results suggest that emphasizing $\beta$, the weight on the electrodogram loss $L_{\rm Elec}$, to 1 does not necessarily improve speech intelligibility. On the other hand, a smaller $\beta$ value of 0.25 limits the refinement of the electrodogram. Among the tested values, $\beta$ = 0.5 results in the highest STOI and ESTOI scores. Although $\beta=1$ achieves a slightly higher NCM score (0.5217) compared to $\beta=0.5$ (0.5211), the difference is considerably small. Therefore, the $\beta$ value of 0.5 is selected for the subsequent experiment.
Table 3: Objective evaluation of various CI implementations under noisy conditions